Problem 14-15 Early extinguishment; effective interest [LO14-5) The long-term liability section of Twin Digital Corporations balance sheet as of December 31, 2017, included 12% bonds having a face amount of $35 million and a remaining discount of $1 million. Disclosure notes indicate the bonds were issued to yield 14%. Interest expense is recorded at the effective interest rate and paid on January 1 and July 1 of each year. On July 1, 2018, Twin Digital retired the bonds at 104 ($36.4 million) before their scheduled maturity Required 1.& 2. Prepare the necessary journal entries for Twin Digital on July 1, 2018. (If no entry is required for a transaction/event, select No journal entry required in the first account field. Enter your answers in whole dollar.) View transaction list Journal entry worksheet Record the redemption of the bonds. Date July 01, 2018
Answer #1
July-1-18 Interest expense 2,450,000 =35000000*14%/2
          Discount on bonds payable 350,000
          Cash 2,100,000 =35000000*12%/2
July-1-18 Bonds payable 35,000,000
Loss on early extinguishment 2,050,000
          Discount on bonds payable 650000 =1000000-350000
Cash 36,400,000